OpenStudy (anonymous):

Describe the vector as an ordered pair. Round the coordinates to the nearest tenth.

OpenStudy (shaik0124):

unable to view ur question it is asking to login u can take scrren shot of that

OpenStudy (anonymous):

How?

OpenStudy (shaik0124):

r u using windows

OpenStudy (anonymous):

YEah

OpenStudy (shaik0124):

windows7?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

@shaik0124

OpenStudy (shaik0124):

@jim_thompson5910

OpenStudy (anonymous):

I said yes

OpenStudy (shaik0124):

JUST A SEC

OpenStudy (agent0smith):

Use the snipping tool in the start menu - just search it, I think windows 7 has it. It'll let you take a screenshot.

OpenStudy (shaik0124):

ALT+PRTSCR

OpenStudy (shaik0124):

will also take a sccreen shot

OpenStudy (anonymous):

Wrong pic,here is the right one

OpenStudy (agent0smith):

Draw it as a triangle like so:  x will be the horizontal component of the vector. y will be the vertical component.

OpenStudy (anonymous):

I need to know the vectors of an ordered pair.

OpenStudy (agent0smith):

This is how you find it. cosine is adj/hyp so: \[\cos 69 = \frac{ x }{ 54 }\] sine is opp/adj so \[\sin 69 = \frac{ y }{ 54 }\] The vector as an ordered pair will be (x, y)

OpenStudy (shaik0124):

x=54*c0s69 y=54 *sin 69 now calculate x and y

OpenStudy (anonymous):

53.643080505=x -6.19837994429=y

OpenStudy (anonymous):

What would the nearest tenth be though?

OpenStudy (anonymous):

@shaik0124

OpenStudy (shaik0124):

solve in similar method as done earler

OpenStudy (agent0smith):

I think you made some mistake in your calculations. x should be smaller than 53.64, and y should be positive, not negative... I found your mistake - your calculator is in radians mode, make sure you change it to degrees.

OpenStudy (shaik0124):

x will be 19.35 in degrees

OpenStudy (anonymous):

50.4133430308=y 19.3518692754=x

OpenStudy (anonymous):

Round to nearest tenth

OpenStudy (shaik0124):

19.4

OpenStudy (agent0smith):

x= 19.351 is closer to 19.4 than 19.3, so you round to 19.4. For y, is 50.41 closer to 50.4, or 50.5? Then your ordered pair is (x, y)